OpenStudy (anonymous):

Andre drove part of a 260 km trip at 80kph and the rest at 100kph. Find the distance he traveled at 80kph if his total time was 2hr 30min

OpenStudy (anonymous):

What do i do after that?

OpenStudy (anonymous):

ahh i see, so i would have to find a common denominator right?

OpenStudy (anonymous):

100t + 80 ( 2.5 - t) = 260 Can you continue from here?

OpenStudy (anonymous):

Yeah thanks! :)

OpenStudy (anonymous):

Would you like to show your work so we can double check ?

OpenStudy (anonymous):

100t + 20 - 8t = 260 100t - 8t = 240 92t = 240 t = 2.61 Is that right?

OpenStudy (anonymous):

How do you distribute 80 into the bracket? 80 ( 2.5 - t) = ....?

OpenStudy (anonymous):

ohh, i thought that was an 8..not an 80

OpenStudy (anonymous):

So could you fix your error now?

OpenStudy (anonymous):

Yes 100t + 200 - 8t =260

OpenStudy (anonymous):

-80 i mean

OpenStudy (anonymous):

Yes, continue please!

OpenStudy (anonymous):

100t + 200 - 80t = 260 100t - 80t = 60 20t = 60 t = 3

OpenStudy (anonymous):

Yes, don't forget the unit t = 3 hrs, though!

OpenStudy (anonymous):

Now find the distance he travels with 80 km/h?

OpenStudy (anonymous):

Right, and now i just plug it in

OpenStudy (anonymous):

is it d = v times t which would be d = 80 x 3 = 240?

OpenStudy (anonymous):

Let me double check!

OpenStudy (anonymous):

All right

OpenStudy (anonymous):

Are you sure the number in your post is correct?

OpenStudy (anonymous):

Since the total of hours is 2.5 hrs, it's impossible that t = 3 hrs

OpenStudy (anonymous):

Yeah im sure..that's what i was thinking..

OpenStudy (anonymous):

Is there no solution?

OpenStudy (anonymous):

If you're sure that your post is correct, then bring this solution to your teacher, point it out that the logic of question is questionable !

OpenStudy (anonymous):

Do you have any question? BTW, you should learn careful when copying down the number and practice your calculating skill more!

OpenStudy (anonymous):

@santosrubi_5 By any chance that you copy the number 260 instead of 230? If so: 20t = 30 -> t = 1.5 hrs => d = 80 ( 2.5 - 1.5) = 80 Km